# Aligning Large Language Models to Follow Instructions and Hallucinate Less via Effective Data Filtering

### v1

@@ 5 Conclusion
In this paper, we introduce NOVA, a novel framework designed to identify high-quality data that aligns well with the LLM’s learned knowledge to reduce hallucination. Our proposed method includes Internal Consistency Probing and Semantic Equivalence Identification, which are designed to separately measure the LLM’s understanding of the given instruction and target response. In this way, we can measure the familiarity of the LLM with the instruction data and prevent the model from being trained on unfamiliar data, thereby reducing hallucinations. Finally, we introduce an expert-aligned reward model, considering characteristics beyond just familiarity to enhance data quality. By considering data quality and avoiding unfamiliar data, we can use the selected data to effectively align LLMs to follow instructions and hallucinate less in the instruction tuning stage. Extensive experiments and analysis show the effectiveness of NOVA.
@@ Appendix D Transferability Study
To verify the transferability of the NOVA method, we conducted experiments on different foundation models using the Alpaca instruction dataset shown in Table 7 and Table 8. We select LLaMA (Touvron et al., 2023) and Qwen-2 (Yang et al., 2024) at the 7B size as the new base models. We aim to gain deeper insights into the applicability of the NOVA method across different frameworks, providing a reference for further research and applications. We find that the NOVA method is also applicable to other models, showing strong transferability and robustness to other models and further research. Compared to other baselines, NOVA significantly reduces hallucinations and keeps a strong ability to follow instructions.
@@ Appendix E Design Exploration
The Design of NLI Model  We further explore the effects of the NLI model on the final performance of NOVA. We first attempt to analyze the effect of the size of the model on the final results. Specifically, we introduce DeBERTa-base-mnli [^4] https://huggingface.co/microsoft/deberta-base-mnli, DeBERTa-xlarge-mnli [^5] https://huggingface.co/microsoft/deberta-xlarge-mnli and DeBERTA-xxlarge-mnli [^6] https://huggingface.co/microsoft/deberta-v2-xxlarge-mnli. As shown in 9, we can find that increasing the size of the NLI model can provide some improvement in the final result, especially when changing the DeBERTa-base-mnli to DeBERTa-large-mnli. However, continuing to increase the model parameters did not have a significant impact on the final performance. Therefore, in order to balance the performance and the inference time of NLI models, we select the DeBERTa-large-mnli to report the final results in our paper. Meanwhile, we further explore whether we use the advanced LLMs (e.g., GPT-4o and GPT-3.5-Turbo) to directly identify the semantic equivalence and get the correct semantic clusters. Specifically, we use the prompt shown in Figure 6 to test the generated responses and the target response by querying the advanced LLMs to identify semantic equivalence. We use the same method as SEI, utilizing the outputs of advanced LLMs to derive semantic clusters and calculate the score of $F_{res}(r)$. As shown in Table 9, the direct application of results from advanced LLMs proves effective in identifying semantic equivalence. Nevertheless, using NLI models deliver competitive or superior final performance while avoiding API-related costs. Consequently, employing NLI models to identify semantic equivalence is both efficient and effective, substantiating the efficacy of our designed SEI approach.
@@
The Design of Quality Reward Model  We also explore the effectiveness of the quality reward model. We introduce UltraFeedback (Cui et al., 2024) and sample 100 instructions and their corresponding responses as the test set. Specifically, for each instruction, we randomly select 2 responses and determine the ranking between the responses based on their labeled scores of instruction-following, honesty, truthfulness, and helpfulness. Only if all four scores are higher will the response be considered a high-quality response. Meanwhile, we involve two Ph.D. students to conduct the human evaluation to ensure the correctness of the response ranking of each sample. Afterwards, we take the instructions and the responses as inputs to each model, and let the model determine the ranking between the responses and calculate the accuracy of the model’s prediction of the ranking. We compare our used Quality Reward Model with GPT-3.5-Turbo-0125 and GPT-4o-2024-05-13. We use the same prompt for each model as Ge et al. (2024). As shown in Table 10, our reward model achieves better performance, showing the effectiveness of our method. Despite GPT-4o’s strong alignment with human preferences in most general tasks, our reward model trained on the expert-revised preference dataset can perform better, highlighting the subtle gap between expert preferences and advanced GPT-4o preferences.
@@ Appendix F Human Evaluation
During the human evaluation, the participants follow the principles in Figure 7 to make the decision. For each comparison, three options are given (Ours Wins, Tie, and Vanilla Fine-tuning Wins) and the majority voting determines the final result. We invite three Ph.D. students to compare the responses generated by the models. Before participants begin to make judgments, we describe the principles of our design in detail and ensure that each participant correctly understands the principals. If the final result can not be determined by the majority voting, we will make the discussion among the participants and vote on the result again.
@@
1. Factuality: Whether the biography provides relatively more factual statements over the non-factual statements? 2. Helpfulness: Whether the biography provides useful information? 3. Relevance: Whether the statements contained in the biography is relevant to the provided people enitity? 4. Naturalness: Whether the biography sounds natural and fluent? Finally, please make a decision among 3 opinions, including Win, Tie, and Loss.

### v2

@@ 5 Conclusion
In this paper, we introduce NOVA, a novel framework designed to identify high-quality data that aligns well with the LLM’s learned knowledge to reduce hallucination. NOVA includes Internal Consistency Probing and Semantic Equivalence Identification, which are designed to separately measure the LLM’s familiarity with the given instruction and target response, then prevent the model from being trained on unfamiliar data, thereby reducing hallucinations. Lastly, we introduce an expert-aligned reward model, considering characteristics beyond just familiarity to enhance data quality. By considering data quality and avoiding unfamiliar data, we can use the selected data to effectively align LLMs to follow instructions and hallucinate less in the instruction tuning stage. Experiments and analysis show the effectiveness of NOVA.
@@ Appendix D Transferability Study
To verify the transferability of the NOVA method, we conducted experiments on different foundation models using the Alpaca instruction dataset shown in Table 7 and Table 8. We select LLaMA (Touvron et al., 2023) and Qwen-2 (Yang et al., 2024) at the 7B size as the new base models. We aim to gain deeper insights into the applicability of the NOVA method across different models, providing a reference for further research and applications. We find that the NOVA method is also applicable to other models, showing strong transferability and robustness to other models and further research. Compared to other baselines, NOVA significantly reduces hallucinations and keeps a strong ability to follow instructions.
@@ Appendix E Design Exploration
The Design of NLI Model  We further explore the effects of the NLI model on the final performance of NOVA. We first attempt to analyze the effect of the size of the model on the final results. Specifically, we introduce DeBERTa-base-mnli [^4] https://huggingface.co/microsoft/deberta-base-mnli, DeBERTa-xlarge-mnli [^5] https://huggingface.co/microsoft/deberta-xlarge-mnli and DeBERTA-xxlarge-mnli [^6] https://huggingface.co/microsoft/deberta-v2-xxlarge-mnli. As shown in Table 9, we can find that increasing the size of the NLI model can provide some improvement in the final result, especially when changing the DeBERTa-base-mnli to DeBERTa-large-mnli. However, continuing to increase the model parameters did not have a significant impact on the final performance. Therefore, in order to balance the performance and the inference time of NLI models, we select the DeBERTa-large-mnli to report the final results in our paper. Meanwhile, we further explore whether we use the advanced LLMs (e.g., GPT-4o and GPT-3.5-Turbo) to directly identify the semantic equivalence and get the correct semantic clusters. Specifically, we use the prompt shown in Figure 6 to test the generated responses and the target response by querying the advanced LLMs to identify semantic equivalence. We use the same method as SEI, utilizing the outputs of advanced LLMs to derive semantic clusters and calculate the score of $F_{res}(r)$. As shown in Table 9, the direct application of results from advanced LLMs proves effective in identifying semantic equivalence. Nevertheless, using NLI models delivers competitive or superior final performance while avoiding API-related costs. Consequently, employing NLI models to identify semantic equivalence is both efficient and effective, substantiating the efficacy of our designed SEI approach.
@@
The Design of Quality Reward Model  We also explore the effectiveness of the quality reward model. We introduce UltraFeedback (Cui et al., 2024) and sample 100 instructions and their corresponding responses as the test set (we find that most of the selected data are in English, but some of the selected instruction types are translation tasks, so a few data contain Chinese responses). Specifically, for each instruction, we randomly select 2 responses and determine the ranking between the responses based on their labeled scores of instruction-following, honesty, truthfulness, and helpfulness. Only if all four scores are higher will the response be considered a high-quality response. Meanwhile, we involve two Ph.D. students to conduct the human evaluation to ensure the correctness of the response ranking of each sample. Afterwards, we take the instructions and the responses as inputs to each model, and let the model determine the ranking between the responses and calculate the accuracy of the model’s prediction of the ranking. We compare our used Quality Reward Model with GPT-3.5-Turbo-0125 and GPT-4o-2024-05-13. We use the same prompt for each model as Ge et al. (2024). As shown in Table 10, our reward model achieves better performance, showing the effectiveness of our method. Despite GPT-4o’s strong alignment with human preferences in most general tasks, our reward model trained on the expert-revised preference dataset can perform better, highlighting the subtle gap between expert preferences and advanced GPT-4o preferences.
@@ Appendix F Human Evaluation
During the human evaluation, the participants follow the principles in Figure 7 to make the decision. For each comparison, three options are given (Ours Wins, Tie, and Vanilla Fine-tuning Wins) and the majority voting determines the final result. We invite three Ph.D. students to compare the responses generated by the models. Before participants begin to make judgments, we describe the principles of our design in detail and ensure that each participant correctly understands the principles. If the final result can not be determined by the majority voting, we will make the discussion among the participants and vote on the result again.
@@
1. Factuality: Whether the biography provides relatively more factual statements over the non-factual statements? 2. Helpfulness: Whether the biography provides useful information? 3. Relevance: Whether the statements contained in the biography relevant to the provided people entity? 4. Naturalness: Whether the biography sound natural and fluent? Finally, please make a decision among 3 opinions, including Win, Tie, and Loss.
@@
Appendix G Case Study for Selected Samples
To evaluate our proposed NOVA qualitatively, we also select some instruction samples from the Alpaca dataset for case studies as shown in Figure 8. Firstly, we can find that simply using $R_{familiarity}$ in Eq. (10) can effectively identify the simple and straightforward instruction samples that align well with LLM’s knowledge. On the contrary, the sample ranked last according to $R_{familiarity}$ contains the open-ended instruction and the very subjective target response. Meanwhile, further using $R_{final}$ in Eq. (10) that considers characteristics beyond just familiarity $R_{familiarity}$ by introducing the quality reward model further enhances data quality, e.g., the complexity of instructions. We can also observe that the sample ranked last according to $R_{final}$ also contains the open-ended instruction and the detailed target response, which may introduce familiar knowledge to LLMs and further promote hallucinations.
[…]
Figure 8: Case study for selected samples.